The problem is 6x(squared)+49x-45=0
I'm not sure if I'm solving this correctly, so could someone please help me step by step by either solving by factoring or solving by using quadratic formula. Thanks!

Answer by stanbon(75887)

6x(squared)+49x-45=0
----
6x^2 + 49x -45 = 0
(x+9)(6x-5) = 0
---
x = 9 or x = 5/6
=====================
Cheers,
Stan H.
============